OpenStudy (anonymous):

Can you help me solve a fraction cross multiplying kind of problem for Chemistry work?

OpenStudy (anonymous):

For tare panda, anything.

OpenStudy (anonymous):

\[\frac{ x}{ 30.07} \times \frac{ 7 }{2 } + \frac{ 9.780 - x }{ 44 } \times \frac{ 5 }{ 1 } = 1.120\]

OpenStudy (anonymous):

Thank you, please help me I really need this for my homework. I tried different ways but I'm still stuck

OpenStudy (anonymous):

Prolly the prettiest way is to make everything on the left have the same denominator \[\frac{7x}{60.14} + \frac{5(9.780-x)}{44} = 1.120 \\ \ \\ \ \\ \frac{(44)(7x)}{(44)(60.14)} +\frac{(60.14)(5(9.780)-5x)}{(60.14)44} = 1.120\] does that make sense?

OpenStudy (anonymous):

Yes so far

OpenStudy (anonymous):

\[ \frac{(44)(7x) + (60.14)(5(9.780)-5x}{(44)(60.14)}= 1.120 \] \[ (44)(7x) - 60.14(5x) + (5)(60.14)(9.780)= (44)(60.14)(1.120) \] \[ x( \ 7(44) - 5(60.14))= (44)(60.14)(1.120) - (5)(60.14)(9.780)\] Then calculate and voila! ^_^

OpenStudy (anonymous):

I'm on it! I'll let you know what I got

OpenStudy (anonymous):

Ok I got it thank you so much!

OpenStudy (anonymous):

^_^ welcome